# Reusing Historical Trajectories in Natural Policy Gradient via Importance Sampling: Convergence and Convergence Rate

### v1

@@ 3 Convergence Analysis
In this section, we analyze the convergence behavior of RNPG by the ordinary differential equation (ODE) method. Throughout the rest of the paper, we only reuse historical trajectories on the gradient estimator $\widehat{\nabla\eta}(\theta_{n})$, while we use trajectories generated by the current policy to estimate the inverse FIM $\widehat{F}^{-1}(\theta_{n})=\widetilde{F}^{-1}(\theta_{n})=\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{n}^{i},\theta_{n})\right)^{-1}$. One of the reasons for such procedure is the high computational cost for estimating the inverse FIM with reusing historical trajectories. We will numerically show that, reusing historical trajectories in estimating inverse FIM only slightly improves the performance with much more significantly increased computational time.
@@ 3.2 Asymptotic Convergence by the ODE Method
where $\bar{F}^{-1}(\theta)=\mathbb{E}\left[\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{i},\theta)\right)^{-1}\right]$, $\xi_{1},\ldots,\xi_{B}$ are $i.i.d.$ samples from the occupancy measure $d^{\pi_{\theta}}$ and $\mathcal{C}(\theta)$ is the Clarke’s normal cone to $\Theta$, $z$ is the projection term, which is the minimum force needed to keep the trajectory of the ODE (6) from leaving the parameter space $\Theta$. The solution trajectory $\{\theta_{n}\}$ in Algorithm 1 also converges w.p.1 to the limit set of the ODE (6).
Note that the positive definiteness of the matrix $\bar{F}^{-1}(\theta)$ implies that the linear system $\bar{F}^{-1}(\theta)\nabla\eta(\theta)=0$ has unique solution $\nabla\eta(\theta)=0$. Therefore, Theorem 1 indicates that the solution trajectory $\{\theta_{n}\}$ in Algorithm 1 converges w.p.1 to a limiting point $\bar{\theta}$ such that $\nabla\eta(\bar{\theta})=0$. Before the formal proof of Theorem 1, we first give a high-level proof outline. Note that in the update (5), we can decompose the natural gradient estimation into three components: the true natural gradient, the noise caused by the simulation error, and the bias caused by reusing historical trajectories. We then separately analyze the noise and bias effects on the estimation of FIM and gradient, and show the noise and bias terms are asymptotically negligible.
@@
Definition 1 (Zero asymptotic rate of change).
@@
where $\delta M_{n}$ is the noise term caused by the simulation error in the gradient estimator, $\zeta_{n}$ is the bias term caused by reusing historical trajectories in gradient estimator, $\delta F_{n}$ is the noise term caused by the simulation error in the FIM estimator. By Lemma 1, the continuous-time interpolations of $\delta M_{n}$ and $\delta F_{n}$ have zero asymptotic rate of change. By Corollary 1, the continuous-time interpolation of $\zeta_{n}$ has zero asymptotic rate of change. Therefore, the limit ODE is determined by the natural gradient $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ and the projection. By Theorem 6.6.1 in [26], the solution trajectory $\{\theta_{n}\}$ in Algorithm 1 also converges w.p.1 to the limit set of the ODE (6). ∎
@@ 4 Characterization of Asymptotic Convergence Rate
Condition (A.2.1) is easily satisfied using the argument in Section 3 under Assumption 1. (A.2.2) further implies $\Sigma_{1}(\theta):=\bar{F}^{-1}(\theta)\Sigma_{\eta}(\theta)(\bar{F}^{-1}(\theta))^{T}$ is continuous in $\theta$. We also define
@@
(A.2.3) requires every eigenvalue of the matrix $\mathcal{G}$ has strictly negative real part, for the sake of asymptotic stability of the differential equation (6).
@@
where sequence $\{\theta_{n}\}$ is obtained by the RNPG algorithm. Let $E_{n}(t)$ denote the piecewise constant right continuous interpolation of $\{E_{m}\}_{m\geq n}$. The following theorem shows the convergence of the normalized error process.
Theorem 2.
@@
where the covariance matrix $\widehat{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$, and $\Sigma_{2}(\bar{\theta})=\Sigma^{\prime}_{2}(\bar{\theta})-\Sigma_{1}(\bar{\theta})$.
@@
where normal-$\Rightarrow$ means convergence in distribution and ${\Sigma_{\infty}}$ has the following expression:
@@
Theorem 2 indicates that the normalized error $E_{n}$ given by the algorithm will asymptotically follow a multivariate normal distribution, whose covariance matrix $\tilde{\Sigma}$ is characterized by $\widehat{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$. The first term characterizes the variation caused by the random samples for the gradient estimator, and the second term $\Sigma_{2}$ reflects the joint impact of random samples for both gradient and inverse FIM estimators. Reusing past samples in the previous $K-1$ iterations helps reduce the covariance caused by the joint impact of random samples for both gradient and inverse FIM by an order of $O(\frac{1}{K})$. We refer the readers to Section 5.4 for the verification of the asymptotic normality on a linear quadratic control (LQC, see [30]) problem.
@@
For ease of notations, denote $\widehat{F}_{n}^{-1}=\widehat{F}^{-1}(\theta_{n})$, $\widehat{\nabla\eta}_{n}=\widehat{\nabla\eta}(\theta_{n})$, $\mathbb{E}_{n}[\cdot]=\mathbb{E}[\cdot|\boldsymbol{e}_{n}(\theta)=\boldsymbol{e}_{n},\theta]$. Note that $\{\delta G_{n}\}$ is a Markov difference sequence, which contains the noise introduced by the new samples, and $g_{n}$ is the conditional mean conditioned on the reused samples. The asymptotic variance of $\{E_{n}\}$ is determined by the asymptotic behavior of the estimator of the natural policy gradient, which can be further characterized through analysis of the two terms: the Markov noise $\delta G_{n}$ and the conditional mean $g_{n}$. We first characterize the asymptotic behavior of $\delta G_{n}$ in the following Lemma 4-6.
Lemma 4.
@@
where the covariance matrix $\widetilde{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{B}\Sigma_{2}(\bar{\theta})$.
@@
4.1 Non-asymptotic Behavior with Constant Step Size
In this subsection, we discuss the non-asymptotic behavior of reusing historical trajectories in the gradient estimation and demonstrate the benefit of reusing with a constant step size. For ease of notation, let $G(\xi,\theta_{n},\theta_{m})=\omega(\xi,\theta_{n}|\theta_{m})G(\xi,\theta_{n})$. The next theorem shows that historical trajectories also helps reduce the variance of the gradient estimation error around the local optima.
Theorem 3.
Under Assumption 1, there exists a constant $C>0$ such that $\omega(\xi,\theta_{n}|\theta_{m})$, $|\eta(\theta)|$, and $\left\|G(\xi,\theta_{n},\theta_{m})\right\|$ are all upper bounded by $C$. For any $\delta>0$ and $n>0$, we have with probability at least $1-\delta$
$\left\|\nabla\eta(\theta^{*})-\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}G(\xi_{m}^{i},\theta^{*},\theta_{m})\right\|_{2}\leq\sqrt{2\beta_{n}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}+\frac{2C}{K}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta},$ (11)
where $\theta^{*}$ is a local optimum, $\beta_{n}=\sqrt{\frac{C^{2}}{K^{3}B(B-1)}\log\frac{\pi^{2}n^{2}(d+1)}{3\delta}}+\frac{1}{K^{2}B}\sum_{m=n-K+1}^{n}\operatorname{trace}(\widehat{\operatorname{Var}}_{m})$, and $\widehat{\operatorname{Var}}_{m}$ is the sample covariance matrix of the vanilla gradient estimator at time $m$, i.e.,
$\widehat{\operatorname{Var}}_{m}=\frac{1}{B-1}\sum_{i=1}^{B}\big{(}G(\xi_{m}^{i},\theta^{*},\theta_{m})-\frac{1}{B}\sum_{j=1}^{B}G(\xi_{m}^{j},\theta^{*},\theta_{m})\big{)}\big{(}G(\xi_{m}^{i},\theta^{*},\theta_{m})-\frac{1}{B}\sum_{j=1}^{B}G(\xi_{m}^{j},\theta^{*},\theta_{m})\big{)}^{T}.$
Note that the trace of the sample variance of the gradient estimation $\operatorname{trace}(\widehat{\operatorname{Var}}_{m})$ is upper bounded by $\frac{B}{B-1}C^{2}$, hence the square of the right-hand side of (11) is of order $O(\max\{\frac{1}{KB},\frac{1}{K^{2}}\})$, which is strictly better than the order $O(\frac{1}{B})$ of the VNPG algorithm. Since the variance of the solution $\theta_{n}$ is determined by the variance of the gradient estimator, Theorem 3 implies that with a constant step size, reusing historical trajectories reduces the variance of the solution around the local optima in RNPG to an order of $O(\max\{\frac{1}{KB},\frac{1}{K^{2}}\})$, compared to the order $O(\frac{1}{B})$ of the VNPG algorithm.
@@ 5.1 Experiment Setting and Benchmarks
For both problems, the policy network is a fully-connected two-layer neural network with 32 neurons and Rectified Linear Unit (ReLU) activation function. We use softmax activation function on top of the neural network. The policy parameter is updated by Adam optimizer. We should note that similar performance can be obtained by using SGD optimizer with an appropriate learning rate. The discount factor is $\gamma=0.99$. We report the average reward over the number of iterations for different algorithms. The reward is averaged over $50$ macro replications. The number of trajectories generated in each iteration (i.e., mini-batch size) is $B=4$. We also set $\epsilon$ to 0.001 to prevent the FIM from becoming singular.
@@ 5.2 Experiment I: Convergence Rate on Cartpole and Inverted Pendulum
Figure 5: Time (s) running RNPG over $n=150$ iterations under different reuse sizes run on cartpole.
@@ 5.4 Verification of Asymptotic Normality on an LQC Problem
In this section, we verify the asymptotic normality of the solution obtained by the RNPG algorithm, as shown in Theorem 2. Consider the following one-dimensional LQC problem with discrete time and discounted cost:
$\min_{u=(u_{1},u_{2},\ldots)}\quad$ $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}(x_{t}+u_{t})^{2}\right]$ (12)
$s.t.\quad$ $x_{t+1}=x_{t}+u_{t}+w_{t},$
@@
Refer to caption $B=5,K=1$
Refer to caption $B=5,K=5$
Refer to caption $B=5,K=10$
Refer to caption $B=10,K=5$
@@
Refer to caption $B=5,K=1$
Refer to caption $B=5,K=5$
Refer to caption $B=5,K=10$
Refer to caption $B=10,K=5$

### v2

@@ 3 Convergence Analysis
In this section, we analyze the convergence behavior of RNPG by the ordinary differential equation (ODE) method. Throughout the rest of the paper, we only reuse historical trajectories on the gradient estimator $\widehat{\nabla\eta}(\theta_{n})$, while we use trajectories generated by the current policy to estimate the inverse FIM $\widehat{F}^{-1}(\theta_{n})=\widetilde{F}^{-1}(\theta_{n})=\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{n}^{i},\theta_{n})\right)^{-1}$. Not reusing samples for the inverse FIM helps simplify the analysis, as correlated samples in the inverse FIM estimator are extremely challenging to handle due to the matrix inversion. Also, our numerical results show that reusing historical trajectories in the inverse FIM estimator, compared to reusing in the gradient estimator, yields only a minor performance improvement. For clarity, we denote the reuse size for the gradient estimator by $K$ instead of $K_{1}$.
@@ 3.2 Asymptotic Convergence by the ODE Method
where $\bar{F}^{-1}(\theta)=\mathbb{E}\left[\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{i},\theta)\right)^{-1}\right]$, $\xi_{1},\ldots,\xi_{B}$ are $i.i.d.$ samples from the occupancy measure $d^{\pi_{\theta}}$ and $\mathcal{C}(\theta)$ is the Clarke’s normal cone to $\Theta$, $z$ is the projection term, which is the maximum force needed to keep the trajectory of the ODE (8) from leaving the parameter space $\Theta$. The solution trajectory $\{\theta_{n}\}$ in Algorithm 1 also converges w.p.1 to the limit set of the ODE (8).
When any solution on the boundary of $\Theta$ is not a local maximum, Theorem 1 indicates that the solution trajectory $\{\theta_{n}\}$ in Algorithm 1 satisfies $\nabla\eta(\theta_{n})\rightarrow 0$ w.p.1.. However, this does not imply the global optimality. Whether the algorithm can achieving global optimality depends on the policy parameterization and the specific problem at hand. For instance, in a simple tabular MDP with direct policy parameterization (i.e., $\pi(\cdot|s)$ for all $s\in\mathcal{S}$), any local optimum is also a global optimum, so NPG with sample reuse will converge to this global optimum. However, in more general MDPs with continuous state or action spaces and more complex policy parameterizations, neither NPG with nor without sample reuse is guaranteed to reach the global optimum. In practice, neural networks are often used to parameterize policies due to their powerful function approximation abilities, which allows for a more flexible policy representation but sacrifices the theoretical guarantee of global optimality.
Before the formal proof of Theorem 1, we first give a high-level proof outline. Note that in the update (6), we can decompose the natural gradient estimation into three components: the true natural gradient, the noise caused by the simulation error, and the bias caused by reusing historical trajectories. We then separately analyze the noise and bias effects on the estimation of FIM and gradient, and show the noise and bias terms are asymptotically negligible.
@@
Definition 1 (Zero asymptotic rate of change)
@@
where $\delta M_{n}$ is the noise term caused by the simulation error in the gradient estimator, $\zeta_{n}$ is the bias term caused by reusing historical trajectories in gradient estimator, $\delta F_{n}$ is the noise term caused by the simulation error in the FIM estimator. By Lemma 1, the continuous-time interpolations of $\delta M_{n}$ and $\delta F_{n}$ have zero asymptotic rate of change. By Corollary 1, the continuous-time interpolation of $\zeta_{n}$ has zero asymptotic rate of change. Therefore, the limit ODE is determined by the natural gradient $\bar{F}^{-1}(\theta)\nabla\eta(\theta)$ and the projection. By Theorem 6.6.1 in [14], the solution trajectory $\{\theta_{n}\}$ in Algorithm 1 also converges w.p.1 to the limit set of the ODE (8). $\blacksquare$
@@ 4 Characterization of Asymptotic Convergence Rate
A sufficient condition for (A.2.1) is assuming a unique local maximum of $\eta(\theta)$ that lies in the interior of $\Theta$, as indicated by Theorem 1. (A.2.2) further implies $\Sigma_{1}(\theta):=\bar{F}^{-1}(\theta)\Sigma_{\eta}(\theta)(\bar{F}^{-1}(\theta))^{T}$ is continuous in $\theta$. We also define
@@
(A.2.3) requires every eigenvalue of the matrix $\mathcal{G}$ has strictly negative real part, for the sake of asymptotic stability of the differential equation (8). This can be guaranteed if $\eta(\theta)$ is locally strongly concave in a neighborhood of $\bar{\theta}$, where $\bar{\theta}$ is the limiting point of the sequence $\{\theta_{n}\}$ in (A.2.1) (please see Appendix A.1 for the proof of this statement. (A.2.5) covers a wide range of diminishing step size such as $\alpha_{n}=n^{-\beta}$ with $11/2<\beta<$ 1.
@@
where sequence $\{\theta_{n}\}$ is obtained by the RNPG algorithm. The normalized term $\sqrt{\alpha_{n}}$ represents the order of the convergence of $\theta_{n}$. By studying the convergence of the normalized error, we want to establish an exact characterization of the asymptotic convergence behavior of the error. Let $E_{n}(t)$ denote the piecewise constant right continuous interpolation of $\{E_{m}\}_{m\geq n}$. The following theorem shows the convergence of the normalized error process.
Theorem 2
@@
where $\widehat{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$, and $\Sigma_{2}(\bar{\theta})=\Sigma^{\prime}_{2}(\bar{\theta})-\Sigma_{1}(\bar{\theta})$.
@@
where $\Rightarrow$ means convergence in distribution and ${\Sigma_{\infty}}$ has the following expression:
@@
Theorem 2 suggests that the random error given by the algorithm, $\theta_{n}-\bar{\theta}$, approximately follows a multivariate normal distribution with mean 00 and covariance matrix $\alpha_{n}\Sigma_{\infty}$ when $n$ is large. We refer the readers to Section 5.5 for the verification of the asymptotic normality on a linear quadratic control (LQC, see [3]) problem. With this approximation, we have,
$\theta_{n}-\bar{\theta}\sim\mathcal{N}(0,\alpha_{n}\Sigma_{\infty}),$
where $\operatorname{vec}(\Sigma_{\infty})=-(\mathcal{G}\oplus\mathcal{G})^{-1}\operatorname{vec}\left(\widehat{\Sigma}(\bar{\theta})\right)$ and $\widehat{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$. In the expression of $\widehat{\Sigma}$, the first term $\Sigma_{1}$ characterizes the variation caused by the random samples for the gradient estimator, and the second term $\Sigma_{2}$ reflects the joint impact of random samples for both gradient and inverse FIM estimators. Reusing past samples in the previous $K-1$ iterations helps reduce the covariance caused by the joint impact of random samples for both gradient and inverse FIM by an order of $O(\frac{1}{K})$. However, as $K$ increases further, this benefit diminishes because the asymptotic variance $\widehat{\Sigma}$ becomes dominated by $\frac{1}{B}\Sigma_{1}(\bar{\theta})$. As a result, the convergence rate of $\theta-\bar{\theta}$ can be characterized, when $n$ is large, as $\theta-\bar{\theta}=O(\alpha_{n}^{\frac{1}{2}}\sqrt{\frac{1}{K}+O(1)})$. Notably, a larger $K$ also demands additional computational resources and memory to store historical samples. In practice, the reuse size $K$ can be chosen based on the computational budget constraints or by balancing convergence gains with computational efficiency and memory usage. We will further explore the impact of reuse size numerically in Section 5.3.
@@
For ease of notations, denote $\widehat{F}_{n}^{-1}=\widehat{F}^{-1}(\theta_{n})$, $\widehat{\nabla\eta}_{n}=\widehat{\nabla\eta}(\theta_{n})$, $\mathbb{E}_{n}[\cdot]=\mathbb{E}[\cdot|\boldsymbol{e}_{n}(\theta)=\boldsymbol{e}_{n},\theta]$. Note that $\{\delta G_{n}\}$ is a martingale difference sequence, which contains the noise introduced by the new samples, and $g_{n}$ is the conditional mean conditioned on the reused samples. The asymptotic variance of $\{E_{n}\}$ is determined by the asymptotic behavior of the estimator of the natural policy gradient, which can be further characterized through analysis of the two terms: the Markov noise $\delta G_{n}$ and the conditional mean $g_{n}$. We first characterize the asymptotic behavior of $\delta G_{n}$ in the following Lemma 4-6.
Lemma 4
@@
where $\widetilde{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{B}\Sigma_{2}(\bar{\theta})$ is the covariance matrix of $W^{\prime}(\cdot)$.
@@ 5.1 Experiment Setting and Benchmarks
For both problems, the policy network is a fully-connected two-layer neural network with 32 neurons and Rectified Linear Unit (ReLU) activation function. We use softmax activation function on top of the neural network. The policy parameter is updated by Adam optimizer. We should note that similar performance can be obtained by using SGD optimizer with an appropriate learning rate. The discount factor is $\gamma=0.99$. We report the average reward over the number of iterations for different algorithms. The reward is averaged over $50$ macro replications. The number of trajectories generated in each iteration (i.e., mini-batch size) is $4$. We also set $\epsilon$ to 0.001 to prevent the FIM from becoming singular.
@@ 5.2 Experiment I: Convergence Rate on Cartpole and Inverted Pendulum
Figure 5: Time (s) running RNPG over $n=100$ iterations under different reuse sizes run on cartpole.
@@ 5.3 Experiment II: Empirical Study on Reuse Size
5.4 Reuse of Samples in Proximal Policy Optimization (RPPO)
In previous sections, we evaluated the performance of policy gradient and natural policy gradient methods with and without reusing samples from past iterations. Here, we investigate the impact of sample reuse in Proximal Policy Optimization (PPO), a state-of-the-art policy optimization method. PPO updates the policy by maximizing the following surrogate function (see [26]):
$L^{CLIP}(\theta)=\frac{1}{B}\sum_{i=1}^{B}\min\left(r^{i}_{n}(\theta){A}^{i}_{n},\operatorname{clip}\left(r^{i}_{n}(\theta),1-\epsilon,1+\epsilon\right){A}^{i}_{n}\right),$
where $r^{i}_{n}(\theta)=\frac{\pi_{\theta}(a_{n}^{i}|s_{n}^{i})}{\pi_{\theta_{n}}(a_{n}^{i}|s_{n}^{i})}$ is the likelihood ratio, $A_{n}^{i}$ is the advantage function, and $\operatorname{clip}(x,a,b)$ is the clipping function.
To reuse past samples, we define the following surrogate function, incorporating samples from the previous $K-1$ iterations:
$L^{CLIP}_{R}(\theta)=\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}\min\left(r^{i}_{m}(\theta){A}^{i}_{m},\operatorname{clip}\left(r^{i}_{m}(\theta),1-\epsilon,1+\epsilon\right){A}^{i}_{m}\right).$
The updated solution $\theta_{n+1}$ maximizes $L^{CLIP}_{R}$. When $K=1$, $L^{CLIP}_{R}$ reduces to the original $L^{CLIP}$.
We tested PPO and its sample-reuse variant (RPPO) on the same Cartpole environment introduced in 5.1. The clip ratio $\epsilon$ is set to $0.2$. Figure 7 shows the mean and standard error of rewards over 150 iterations for both PPO and RPPO with reuse sizes $K=2,5$, averaged over 50 replications. Results indicate that RPPO achieves higher average rewards and lower standard errors with larger reuse sizes, demonstrating the benefit of sample reuse.
Figure 7: Mean (Figure 7(a)) and standard error (Figure 7(b)) of the reward over $n=150$ iterations for RPPO under reuse sizes $K=1,2,5$ run on cartpole.
@@ 5.5 Verification of Asymptotic Normality on an LQC Problem
In this section, we numerically verify the asymptotic normality of the random error in the solution obtained by RNPG, as shown in Theorem 2. Consider the following one-dimensional Linear Quadratic Control (LQC) problem with discrete time and discounted cost:
$\min_{u=(u_{1},u_{2},\ldots)}$ $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}(x_{t}+u_{t})^{2}\right]$ (13)
$s.t.$ $x_{t+1}=x_{t}+u_{t}+w_{t},$
@@
$B=5,K=1$
$B=5,K=5$
$B=5,K=10$
$B=10,K=5$
@@
$B=5,K=1$
$B=5,K=5$
$B=5,K=10$
$B=10,K=5$